OpenStudy (anonymous):

If x=2pi/2009, then cosxcos2xcos3x....cos1004x = ?

OpenStudy (anonymous):

@ganeshie8 @eliassaab

OpenStudy (anonymous):

I did it numerically, I found the quantity equal \[ 5.8329*10^{-303}\approx 0 \]

OpenStudy (anonymous):

Yes, you are going to have some values of cos@ very near zero, for example in the vicinity of 250 x=(500/2009) pi = pi/4 approx.

OpenStudy (anonymous):

But\[ \cos\left ( \frac \pi 4 \right)=\frac {\sqrt 2}{2} \]

OpenStudy (anonymous):

Yes, but multiplication by near-zero values wipes all else out.

OpenStudy (anonymous):

You want to consider \[ 500 x \approx \frac \pi 2 \]

OpenStudy (anonymous):

\[ \cos\left ( \frac \pi 2 \right)=0 \]

OpenStudy (anonymous):

I am not getting equations from you but formatting instructions, hard to follow. As your numerical calculations confirmed, those near-zero cosines wipe out everything else.

OpenStudy (anonymous):

Also all of theses cosines are < 1, the product of 1004 of them must be very small.